Walderland is a fictitious country where the local currency is called waldis. At present, the  currency held by the public is 500,000 waldis, bank reserves are 500,000 waldis, and the required  reserve-deposit ratio is 0.10. Assume throughout the problem that the public does not wish to change  the amount of currency they hold. The central bank in Walderland conducts an open market sale of  200,000 waldis. What is the new money supply?


1
Expert's answer
2021-05-03T10:53:13-0400

The money supply is given by:

"M=m\times MB"

Where m is the money multiplier and MB is the monetary base.

The monetary base is the sum of the total reserves and the currency in circulation.

"MB=C+R"

The money multiplier is equal to:

"m=\dfrac{1}{r}"

Where r is the required reserve ratio.

If the reserve requirement is 0.10, then the money multiplier is equal to:

"m=\dfrac{1}{0.10}=10"

Therefore, the money multiplier supply becomes:

"M=10\times MB"

If the currency in circulation is 500,000 and the total reserves are 500,000, then the monetary base is:

"MB=500,000+500,000=1,000,000"

If there is an open market conduct of 200,000, the monetary base will increase by 200,000 and the new monetary base is:

"MB=1,200,000"

Thereforem the new money supply is:

"M=10\times 1,200,000\\[0.4cm]
M=\boxed{\color{red}{12,000,000}}"